Fermi National Laboratory
 

Fermilab Shutdown Report for July 10, 2009 - July 17, 2009

 

Fermilab Shutdown Report
7/10/09 – 7/17/09

LINAC

Shutdown Week 5 – On Schedule

  • DT20 work is complete, Tank Five is closed and under vacuum
  • Pump down in progress of LE tanks 4 & 5 and Transition section
  • On-going maintenance of LE and HE Stations
  • Going to controlled access next week
  • Expected return of NTF operations on week of 27 th of July

Booster

Shutdown Weeks 5 – On Schedule

  • Correction element upgrade ~83% with magnet removal/installation
    • 30 out of 36 magnets installed. 30 Surveyed and under vacuum. 19 magnets terminated in tunnel
    • All of Booster is currently under vacuum
    • Old correctors in period 3-6 (6 magnets) Will be pulled starting Monday 7/20
    • Corrector Wire pulls 100% complete
    • Other PS work 80% completed
  • East Gallery LCW Cooling pipe upgrade 100% completed. Still need to fill system. Tested good with nitrogen pressure test
  • CUB work on Booster 95 system has started

GMPS Upgrade Status – New East Yard Brentford 3 & 4

  • GMPS upgrade ~90% complete
  • Making Wiring connections on switchgear and Transformers
  • Work on GMPS 1 & 2 also in progress

Sump and ICW Leaks Repaired

  • Problem with shut-off valve to Booster Pond. It will not open and needs to be replaced. Valve is in Booster Ring Road.
  • Will lower pond level Monday and Diver will install plug in Pond discharge line.
  • Excavation to start Tuesday.
  • ICW leak found and fixed in Booster East Lot. Disintegrated bolts found on indicator valve stem.

Antiproton Source

Accumulator 4-8GHz DP Kicker Upgrade: ~70% complete

  • The A50-A60 section is under vacuum and leak checked. Preparation for the A60 bake is being done today. The bake will start on Monday morning and finish on Friday

Debuncher Septum Magnet Replacement: 50% complete

  • Analysis of cracked weld was done and it has been decided to repair the broken magnet. When repairs are complete, it will be placed under vacuum ASAP and pulsed tested under full current for an extended period to watch the vacuum. Analysis of copper debris in aperture shows it was molten when created

IB7 (AP2 dipole magnet) Replacement: 50% complete

  • As of this morning, both the old magnet and the replacement are sitting in the AP0 service building. The plan is to have the replacement in the tunnel by the end of today

Other Work

  • Maintenance work on various RF systems continues.
  • LCW leaks being addressed.
  • Power supply maintenance – LCW hoses, filter capacitors, etc
  • Target SEM work continues

Main Injector

MI-10 area: On Schedule

  • All cores bored
  • 14 of 18 conduits in place, grouted, starting back fill on 6
  • Building foundation progressing quickly

MI-40 area: On Schedule

  • All cores bored
  • All 12 conduits in place
  • Tunnel top sealed, back filled to >5ft
  • Building construction not yet started

Tunnel Work

MI-8

  • Beam valves relocated
  • Multiwires being assembled/tested/prepped for installation
  • New ceramic MW @ 810, 811, 813, 814 & 830 for E + dp/p measurements. (maybe 851)

MI-10

  • Still have 2 pans mounted to ceiling for construction

MI-40

  • Beam valve and kickers moved
  • Gap Clearing kickers being assembled
  • Kicker stand installation can now continue (tunnel re-opened 7/16)

MI LCW

  • 12 quad locations left to re-hose
  • Aligning/repairing LCW pumps around the ring
  • Replacing/calibrating pots in 3-way valves
  • RF heat exchangers cleaned

MI PS

  • Too much work to list. All maintenance is going well: on schedule

MI eCloud (MI-52)

  • Added eCloud BPMs at 518, 519.
  • Preparing to install new beam valve and eCloud components @521

MI RF

  • Modulators cleaned, strainers cleaned, all work on schedule

Recycler

Recycler

  • Pot RR129 (i.e. MI-10 injection kicker)
    • Not started
      • Expected to start on Wednesday (and be completed by the end of the week)
  • Move 6 magnets ( complete)
  • Vacuum maintenance
    • Not started
  • Power supplies maintenance
    • VDPA1 & VDPA2 maintenance (75%)
      • No work done this week
    • Correctors (not started)
  • HLS installation at MI-10 (possibly)
    • ALARA planning in progress

Electron Cooling

  • Pelletron maintenance (30% complete)
    • Slowed down by Recycler venting
      activities
      • Vacuum issues with manifold
    • Chains, sheaves have been replaced
      • Need running test in air
        • Needs Safety approval
        • Monday or Tuesday
  • Flush Water/Glycol skid (and magnets) (40% complete)
    • All lines, skids and magnets have been flushed
      • Hoses were replaced and manifolds cleaned (ultra-sound)
  • NMR probes replacement (80% complete)
    • Tuning on the probes in the 180 ° bend magnet in progress

Recycler Magnet Motion at MI-40 during excavation

  • Magnets moved up ~8 mm when five feet of soil was removed

Recycler Pump Out

  • Guidelines – the following will begin as soon as the contractors finish their tunnel work
    • As soon as possible
      • Needs to be balanced with danger caused by the remaining work
    • Needs some time pumping to find out if baking is required
      • Estimate 1 week
    • Baking adds 3 weeks to the schedule
    • First protons on September 1 st (according to the current schedule)
      • Pumping down of the Recycler should take place no later than August 1 st

Tevatron

Other Items

  • Un Roll TeV magnet 80% complete
  • As founds on feed cans and pipes at E3
  • E3 feed can replaced interfacing underway
  • Crystal for Vertical collimator damage will have only 1 crystal
  • Repairing HLS systems around Tevatron D0 system up B0 system next week
  • Turbo pump work around the ring
  • Sump pump and dehumidifier maintenance around the ring
  • Conning tower at E3 has loose connection TD will replace Friday
  • Will have day and night shift for vacuum work starting July 27

NuMI

Shutdown Weeks 5

  • Completed Modification of Horn Power Supply for Anti-Neutrino Mode
  • Complete Repair of Horn 2 Water Leak of EVAC Seal between the Horn and Bellows to the Water Tank
    • Practiced Repair Procedure & Reviewed ALARA Plan
    • Moved Horn 2 into Work Cell & Set Up Temp Shield Wall
    • Removed Damaged EVAC Seal on Tuesday
    • Evaluate Changes in Procedure to Install New Seal
    • Installed New EVAC Seal on Wednesday
    • Verified Repair by Smoke Test & Completed Hipot in Work Cell
    • Moved Horn 2 Module back into Chase & Complete Hipot
    • Install Water Lines and Start RAW Skid
  • •Profile Monitor Replacement Completed on PM101 & PM105 & Alignment Complete
    • Installed Fermi Profile Monitors @ Both Locations
  • EPB Dipoles PEEK Installation Continues
    • HV101-1, HV101-2, HV101-6 Installed & Alignment Completed
    • Work Continues on HV101-3, HV101-4, HV101-5
  • Completed Initial Mobilization for Hadron Monitor Replacement
  • Continue Installation of Controls System for Absorber Hall Dehumidification System

Cryogenic System

TeV at LN2 Temperatures

Operations and Maintenance Status

  • Continued support AD/MSD for A3, D3, & E3 repairs; F1 warming up to room temperature
  • FRIG A2 low-pressure heat exchanger helium-air leak repaired successfully.
  • Continued ringwide vacuum pump maintenance, repairs, and pumpdown of various components to assure stable Tevatron LN2 Mode Operations
  • 8 wet engine flywheels replaced, 5/5 wet engine motors replaced, 18/26 wet/dry engine overhauled for TeV/Pbar ops, 4/6 IHI cold compressors replaced
  • Mycom Compr Benshaw starter modifications, 7/8 (TEV) completed
  • Mycom Compr relief valves replaced (2/8)
  • Ongoing numerous FRIG/Compr controls/instrumentation maintenance/repairs.
  • CHL at room temperature & maintenance continuing (equip greasing/oil changes, NRL Niagara aftercooler cleaning-50% done, relief valves inspection/testing, instr. repairs. Liquid nitrogen supply to Tevatron relying on Linde deliveries (2-3 per day) to CHL’s 150,000 gallons of tank storage

Controls

Completed This Week:

  • BSSB Update & MTA Switchbox Installed
  • TLG rebooted
  • AD Firewall maintenance
  • NuMI C204 & PCI installed at MI-60N (MADC #77)
  • Beamssrv1 Storage Upgrade
  • Application & Device Database F/W, OS & Sybase S/W updates
  • Linac Controls Maintenance (IRMs, PLCs)
  • New Pulse Shifter
  • General Maintenance

Next Week:

  • Continued Linac Controls Maintenance & Checkout
  • General Maintenance

Reminder:

  • Networks & TCLK will be down weekend of 7/25 & 7/26

MTA

All shield work has been completed.

FESS

Electrical and High Voltage

  • MSS – energized on 7/15 [not on line]. All work completed
  • MSS Findings:
  • 345 kv Air Switch guide ball/hardware corroded off most likely due to arcing
  • Insulator [cracked] replaced on T-83 transformer buss duct
  • MI – HV & electrical PMs ~60% completed [transformers, switch gear, panels etc] WIP
  • TeV & Linac Lighting – Relamping completed except CWOs on ballasts & panels WIP
  • Switch Yard – Oil switch replacement to be scheduled when bid process completed [next week?] – 1 week shutdown required

Mechanical

  • AP-0 – Sumps & Fans completed
  • TEV – Pond pump PMs completed. Finding: F-4 30 hp pond pump to be replaced

HVAC

  • MI – Marv Air SB cooling PMs & repairs WIP & dehumidifiers repairs WIP

CUB

  • Chiller # 3 [55LCW] – the Chiller was serviced & the condenser was chemically cleaned & open for inspection with a finding that showed leakage points of 30 tubes between the tubes & the tube sheet under 8psi N2 pressure. Nine tubes also require plugging per EDDY Current testing. Trane Co. is reviewing the findings for action that could mean a near future condenser replacement or tube replacement. This chiller has been in service for 10+years & runs 24/7 servicing LINAC
  • 55LCW [LP4 &5] – VSDs installed & being configured by DDC group
  • Chiller # 4 [Process] – condenser cleaning/de-scaling in progress. To open/inspect/eddy current next week
  • HE9 Pbar heat exchanger – cleaning WIP
  • HE1 booster heat exchanger – cleaning scheduled next week with a follow-up plate inspection
  • CUB shut down weekend of 7/25 & 7/26. NO COOLING/AC WILSON HALL & other service areas
  • Cooling Tower Water piping repair to HE1 [booster] & HE9 [Pbar] completed
  • 95LCW Booster return pipe bellows/expansion sleeve & valve replacements WIP – 7/24

Other Work and Comments

  • BTE sump line - isolation valve replacement [excavation] scheduled for Tues 7/21 which will close off the south entrance to booster ring area road for ~8hrs. The repair at booster east excavation will remain open until the sump line is leak checked. A project has been submitted to FESS Engineering to replace the entire sump discharge line for 2010
  • BTE – ICW leak repaired in parking lot replacing PIV 109
  • ICW water main at roads A1/A2 - directional boring with 12” line installed. To be operational by 7/31
  • Casey’s Pond – shutdown set for 7/25 to inspect sump for ICW distribution
  • ICW flushing followed by Zebra Mussel treatment starts 8/1 thru 8/8

MiniBooNE

  • Cleaning tunnels
  • Working on the horn power supply

Switchyard 120 – Work has not started

Collider Detector at Fermilab

  • The silicon cooling work I going slow
  • The end plugs should be closed up next week

D-Zero – On Schedule

Mechanical

8 GeV multiwire testing

Ceramic tubes for kicker

Cleaning Linac, Booster, and MI tunnels

Instrumentation– On Schedule

  • Have parts for multiwire

Safety

Safety

  • Shutdown ALARA plans are progressing very close to as planned

Interlock Work

  • Single Line Electrical Drawing Verifications
    • 20% complete
  • Main Injector Interlock Repairs and PLC Code Modifications are complete
    • Thanks to Operations for securing the MI tunnel after the PLC code change was complete
  • Tevatron ground fault diagnostics installation
    • 20% complete
  • Misc. Tevatron Interlock repairs
    • 50% complete
  • Safety System Testing
    • Linac and NTF tentatively scheduled for July 23 rd

Shutdown Safety

  • Injuries
    • None to report since last meeting

NuMI Horn 2 Leak Repairs

  • Initially work not going as planed
    • Crews didn’t force the situation, stopped, developed, and practiced a new plan
    • Revised plan successfully repaired the horn with approximately 90 person mrem of dose
    • Great Job to all involved!

Continue using the Take Five principles Before, During, and After work activities to:

  • Avoid Injuries
  • Reduce Environmental Impacts
  • http://www.fnal.gov/takefive/
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/17/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory